OpenStudy (anonymous):

(please help i am stumped on this one)The surface area of the square pyramid is 96 square inches. Each of the triangular faces of the pyramid has an area of 15 square inches with a height of 4 inches.

OpenStudy (anonymous):

whats the question?

OpenStudy (anonymous):

sorry... whats the volume equal?

OpenStudy (anonymous):

The volume of a pyramid with a rectangular base is equal to: Length_of_base * Width_of_base * Height * 1/3 In your case legth = width. And the area of the base is 96 - 4*15 = 36 so the length and with are equal to 6. So 6*6*4*1/3=?

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yeah thanks